The problem is: -3a^2(5ab-3b+7a) I am not sure between two answers, which are:
-15a^3b+9a^2b-21a^3 or -36a^3+6a^2b
Also can you please tell me if these are correct answers:
(3x-12)(x-1)= 3x^2-15x+12
(8x+y)(8x-y)= 64x^2-y
(3x-12)^2= 9y^2-72y+144
Thank you so much for the help!!!

Answer by stanbon(48569)

-3a^2(5ab-3b+7a) I am not sure between two answers, which are:
-15a^3b+9a^2b-21a^3 or -36a^3+6a^2b
The 1st is correct
--------------------------------
Also can you please tell me if these are correct answers:
(3x-12)(x-1)= 3x^2-15x+12
Correct
-------------------------
(8x+y)(8x-y)= 64x^2-y
No: 64x^2-y^2
--------------------------
(3x-12)^2= 9y^2-72y+144
Correct
------------------
Cheers,
Stan H.